OpenStudy (anonymous):

equivalent form of this equation x/12 -7=x/4 options 1. x-7=3x 2. x-74=12x 3. x-84=3x

mathslover (mathslover):

u can write : \[\huge{\frac{x}{12}-7=\frac{x}{4}}\] \[\huge{\frac{x-84}{7}=\frac{x}{4}}\] U can cross multiply .. can u do this?

OpenStudy (anonymous):

i actually don't know how too:/

mathslover (mathslover):

ok cross mu